# Making Interpretable Discoveries from Unstructured Data: A High-Dimensional Multiple Hypothesis Testing Approach

### v1

@@ Abstract
Social scientists are increasingly turning to unstructured datasets to unlock new empirical insights, e.g., estimating causal effects on text outcomes, measuring beliefs from open-ended survey responses. In such settings, unsupervised analysis is often of interest, in that the researcher does not want to pre-specify the objects of measurement or otherwise artificially delimit the space of measurable concepts; they are interested in “discovery.” This paper proposes a general and flexible framework for pursuing discovery from unstructured data in a statistically principled way. The framework leverages recent methods from the literature on machine learning interpretability to map unstructured data points to high-dimensional, sparse, and interpretable “dictionaries” of concepts; computes (test) statistics of these dictionary entries; and then performs selective inference on them using newly developed statistical procedures for high-dimensional exceedance control of the $k$-FWER under arbitrary dependence. The proposed framework has few researcher degrees of freedom, is fully replicable, and is cheap to implement—both in terms of financial cost and researcher time. Applications to recent descriptive and causal analyses of unstructured data in empirical economics are explored. An open source Jupyter notebook is provided for researchers to implement the framework in their own projects.
@@ 1 Introduction
Empowered by recent developments in machine learning and AI, researchers in the social sciences are increasingly leveraging sources of unstructured data—such as text, images, videos, and audio—in quantitative analyses. In economics, interest in unstructured data sources is especially widespread, ranging from using speech recordings from FOMC meetings to better understand monetary policy [Gorodnichenko et al., 2023]; to using videos of start-up pitches to study entrepreneurship and investment [Hu and Ma, 2025]; to using open-ended survey questions to probe the economic behavior and beliefs of individuals [Haaland et al., 2024]; to using mug shots to study judicial decisionmaking [Ludwig and Mullainathan, 2024]; to using visual art or written narratives to infer long-run living standards [Gorin et al., 2025, Lagakos et al., 2025]; to using interview transcripts to better understand the impacts of RCT treatments [Bergman et al., 2024, Krause et al., 2025].
@@
The promise of these unstructured data, fundamentally, is that they offer new opportunities to measure social phenomena that were previously unmeasurable. This includes both measures that a researcher is capable of pre-specifying as well as those that a researcher is not, and instead would like to “discover.” When modern AI/ML methods are brought to bear on these goals, the former falls under the heading of “supervised learning,” whereby the researcher typically makes predictions of their ex-ante known quantity of interest using the unstructured data as features, and the latter falls under the heading of “unsupervised learning,” which is less amenable to a prediction-based framework, and demands that AI/ML methods uncover latent structure in the unstructured data with minimal researcher input. Though new statistical and econometric frameworks have recently become available for interpretable and statistically principled analyses of unstructured data in the supervised learning setting (e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025]), it is an open question as to how best perform interpretable and statistically principled analyses of unstructured data in unsupervised learning settings that emphasize discovery. The framework proposed in this paper provides one possible answer to this open question.
@@
Specifically, this paper proposes a general framework for conducting interpretable and rigorous statistical inference on unstructured data in the following way, starting from an unstructured dataset $\{Z_{i}\}_{i=1}^{n}$:
@@
    Convert unstructured data to dictionary feature vectors: The researcher converts the unstructured data instances $Z_{i}\in\mathcal{Z}$ into high-dimensional, sparse, binary vectors $Y_{i}\in\{0,1\}^{p}$ using recently developed methods in dictionary learning for mechanistic interpretability of large language models (LLMs) and other deep neural networks (DNNs), i.e., $Y_{i}:=\mathtt{Dict}(Z_{i})$. Both theory and empirics corroborate that vector entries $Y_{ij}$ correspond to the presence of a monosemantic feature (i.e., a feature with a single, interpretable meaning) or “concept” $j$ in the given unstructured data instance $i$. There are $p\gg n$ dictionary features, e.g., we assume that $n$ may be on the order of $10^{2}$ or $10^{3}$, and $p$ is on the order of $10^{4}$ or $10^{5}$ (or larger). The breadth and depth of concepts spanned by these dictionary features is vast, as can be gleaned from the literature on dictionary learning for LLMs, and as will be observed concretely in the empirical applications of this paper. Intuitively, the dictionary should contain all possible (monosemantic) concepts that the LLM has learned for the purpose of next token prediction over a massive internet-scale corpus. However, due to sparsity imposed by design and suggested by machine learning theory, only a handful of these dictionary features will be present (“activate”) for any given unstructured data point. [^1] For more discussion of what “monosemanticity” means and its importance for LLM interpretability, see, e.g., the discussion in Bricken et al. [2023].
@@
    Compute test statistics from the dictionary feature vectors: The researcher computes from this newly created dictionary feature indicator dataset $\{Y_{i}\}_{i=1}^{n}$ a vector of $p$ test statistics, $T_{n}\in\mathbb{R}^{p}$. These test statistics are intended to test $p$ hypotheses, one for each dictionary feature, and may be flexibly chosen by the researcher. Especially important and simple choices explored in this paper’s empirical examples include hypotheses testing means and differences-in-means, feature by feature.
@@
    Perform high-dimensional selective inference using the test statistics: Finally, the researcher invokes a high-dimensional selective inference procedure newly developed in this paper to control the $k$-FWER of the selected set of rejected hypotheses (“discoveries”). This procedure is asymptotically valid for small $k$, requires only very mild assumptions on the data generating process (which are met in the applications of interest), and is well-powered in the empirical examples considered. This procedure may also be of independent interest in other high-dimensional selective inference settings.
The output of this framework is therefore a set of interpretable discoveries (rejections of dictionary feature derived tests) based on some set of unstructured data $\{Z_{i}\}_{i=1}^{n}$ with generalized familywise error rate guarantees, which may be relaxed or tightened as the researcher sees fit to permit more or fewer discoveries.
Two notable use cases for this framework explored in the empirical applications of this paper include:
  1. 1.
    Estimating average treatment effects on dictionary features in RCTs, or testing, for treatment indicator $D_{i}$ and fixed probability of treatment $\pi$,
    $H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
  2. 2.
    Estimating the probability of the presence of various concepts (measured by monosemantic dictionary features) in the population from which the unstructured dataset was sampled, or testing
    $H_{0,j}:E[Y_{ij}]=0$
    and then inverting these tests to form a “generalized” confidence set [Romano and Wolf, 2007].
@@
This particular framework for discovery from unstructured data conveys several scientific benefits. The first is that of having few researcher degrees of freedom. In many settings featuring exploratory analysis of unstructured data, there is concern that researchers may cherry pick aspects of their data to measure or cherry pick choices of measurements in order to reverse engineer a specific conclusion. By design, however, the framework proposed by this paper has very little room for motivated data mining, as it requires no “human-in-the-loop” to direct what quantities are ultimately measured. Off-the-shelf, open-source dictionary learning models generate $p$ features spanning a huge number of concepts that are well-cataloged and publicly documented, making it unlikely the researcher needs to define additional concepts to measure ad hoc, and making a failure to report hypotheses for certain features conspicuous. Moreover, prior to analysis, a researcher could simply preregister a choice of dictionary learning model, such as particular sparse autoencoder (SAE) for a particular LLM, limiting the possibility that a researcher applies this framework with many different dictionary learning methods to select on favorable results. The proposed framework is also inexpensive both in terms of financial cost and researcher burden; both empirical examples in this paper were computed on Google Colab notebooks using a single A100 GPU, access to which only costs tens of dollars per month. [^2] This pricing description is valid as of October 2025. Further, this framework is easy to implement, requiring no human-in-the-loop to label unstructured data instances, no additional AI/ML model training, and no additional data sources; the framework is nearly automatic in execution, and takes only minutes to run using the above-mentioned hardware. This automaticity also lends the framework to fast and easy replication, another important scientific benefit.
@@
    The development of novel high-dimensional multiple hypothesis testing procedures for controlling $k$-FWER exceedance, facilitated by extensions of high-dimensional central limit theory.
The remainder of this paper is structured as follows: Section 2 discusses related literature; Section 3 describes the framework in detail; Section 4 applies the framework to two recent papers making use of unstructured data for discovery [Bursztyn et al., 2023, Stantcheva, 2024]; and Section 5 concludes.
@@ 2 Related Literature
Motivated by the capabilities of modern AI/ML methods for learning from unstructured economic data (see, e.g., Mullainathan and Spiess [2017], Gentzkow et al. [2019a], Dell [2025]), new econometric and statistical frameworks have been developed to facilitate principled statistical inference on low-dimensional features (predictions) learned from unstructured datasets under supervision, e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025], Rambachan et al. [2024]. This most recent wave of econometric literature emphasizes nonparametric frameworks compatible with black-box AI models coupled with debiasing methods, as opposed to model-based approaches, e.g., Gentzkow et al. [2019b], Battaglia et al. [2024]. Though the framework presented in this paper is concerned with principled inference on unstructured data without making parametric assumptions, it differs from this existing econometric literature in that its goal is primarily unsupervised discovery, as opposed to supervised detection, using AI/ML models.
Both the proposed framework and the recent econometric framework of Modarressi et al. [2025] support making causal inferences on text, with an emphasis on discovery. Modarressi et al. [2025] uses LLM reasoning as a tool for dimensionality reduction on the space of concepts present in text (i.e., as a low-dimensional hypothesis generation methodology) and incorporates human validation of LLM-selected hypotheses, with estimation implemented via sample splitting; in contrast, the framework proposed in this paper leverages recent advances in interpretability methods for LLMs in order to directly analyze the high-dimensional space of concepts present in text, and uses state-of-the-art autointerpretation methods to describe discovered concepts without further researcher intervention, and without requiring data splitting. As such, these frameworks offer different profiles of scientific benefits and researcher costs for the problem of causal inference on text.
Modarressi et al. [2025] also directly leverages the work of Ludwig et al. [2017] in their pipeline, which is similarly relevant to the proposed framework, in that both emphasize discovery from high-dimensional data, and can likewise both be applied to the problem of causal inference on unstructured data (as it is used in Modarressi et al. [2025]). However, the proposed framework tackles the high-dimensional discovery problem using a novel high-dimensional selective inference procedure that maintains a high degree of interpretability, whereas Ludwig et al. [2017] modifies the target of discovery to improve power, trading off interpretability to do so.
@@
There is an extensive literature on multiple hypothesis testing (or “selective inference” more broadly) in statistics, biostatistics, and economics (see, e.g., Romano et al. [2010] for a review in econometrics). The methods in this literature span asymptotic and finite sample valid frameworks, and low- and high-dimensional settings. In particular, for the problem considered in this paper, the literature on asymptotically valid, high-dimensional selective inference is most relevant, as such methods permit making very few assumptions about the data (c.f., assuming independence or PRDS of the p-values in the case of Benjamini and Hochberg [1995]) and readily incorporate resampling methods that improve power (relative to guarding against worst-case dependence, c.f., Bonferroni corrections or the method of Benjamini and Yekutieli [2001]). These new procedures are built by extending recent work on high-dimensional central limit theory for order statistics [Chernozhuokov et al., 2022, Ding et al., 2025]. The proposed statistical procedures are the first stated of their kind that enable asymptotic high-dimensional exceedance control of the $k$-FWER for test statistics that may be arbitrarily correlated with one another, complementing existing methods for asymptotic high-dimensional control of the FWER [Belloni et al., 2018] and the FDR under weak dependence [Liu and Shao, 2014, Belloni et al., 2018].
@@
There is a growing literature in economics and computer science on “hypothesis generation” at the pre-scientific stage of empirical inquiry [Ludwig and Mullainathan, 2024]. From the perspective of this literature, the proposed framework may be viewed as leveraging dictionary learning as a high-dimensional hypothesis generator, on top of which discovery is performed at the “scientific” stage of inquiry using high-dimensional multiple hypothesis testing procedures. As such, the framework in this paper contributes to the literature on hypothesis generation by providing a statistically principled pathway from generation to evaluation—one which involves little to no researcher discretion about what generated hypotheses should be tested. The recent framework of Movva et al. [2025] uses dictionary learned features (specifically, sparse autoencoder features) explicitly for the purposes of hypothesis generation, further motivating the use of dictionary learning methods as tools for hypothesis generation in the social sciences, both in the present framework and beyond.
@@
There is a large literature in computer science on machine learning interpretability (see, e.g., Doshi-Velez and Kim [2017] for a foundational agenda). The machine learning interpretability methods leveraged by this framework originate from a nascent though highly active literature known as “mechanistic interpretability,” which seeks to develop methods for interpreting LLM behavior via quantitative analyses of model internals (e.g., activations, weights). In particular, the dictionary learning methods for DNN interpretability implemented in this framework were developed in Bricken et al. [2023], Templeton et al. [2024], i.e., sparse autoencoders (SAEs). SAEs are autoencoders attached to the residual streams (or other model internals) of a LLM, with hidden layers being many orders larger in dimension than the residual stream, and which are trained with sparsity-inducing penalization on reconstruction loss. Under the “linear representation hypothesis” and “superposition hypothesis” (see, e.g., Bricken et al. [2023] for more discussion), these sparse autoencoders are thought to act as an overcomplete basis of the space of concepts in text leveraged by a LLM to make next token predictions, encouraging learned features to be monosemantic. There has been much debate about the success of SAEs (e.g., Leask et al. [2025]), though most criticisms still support the notion that SAEs and other dictionary learning methods have a comparative advantage in discovering as opposed to detecting concepts of interest in text [Peng et al., 2025]. Other dictionary learning methods that have been proposed as competitors to SAEs (e.g., transcoders [Paulo et al., 2025]) are similarly compatible with the framework proposed in this paper.
SAEs and other dictionary learning methods generate features, but not natural language feature descriptions. Based on the popularity of these methods, a related literature on “autointerpretability” methods has become active, which seeks to coherently use LLMs to describe the features discovered by dictionary learning methods at scale. Recent important papers in this literature include: Bills et al. [2023], Shaham et al. [2024], Paulo et al. [2024], Rajamanoharan et al. [2024]. The framework proposed in this paper primarily uses insights from Paulo et al. [2024] to create “local interpretations” that are most relevant to the data generating process being considered, and otherwise leverages autointerpetations from Lieberum et al. [2024].
Though the empirical examples considered in this paper handle one of the most popular unstructured date types—text—for which dictionary learning-based interpretability methods are the most well-developed, dictionary learning techniques have been successfully applied to many other modalities, including audio and images [Abdulaal et al., 2024, Bhalla et al., 2024, Fry, 2024, Daujotas, 2024, Pluth et al., 2025].

### v2

@@ Abstract
Social scientists are increasingly turning to unstructured datasets to unlock new empirical insights, e.g., estimating descriptive statistics of or causal effects on quantitative measures derived from text, audio, or video data. In many such settings, unsupervised analysis is of primary interest, in that the researcher does not want to (or cannot) pre-specify all important aspects of the unstructured data to measure; they are interested in “discovery.” This paper proposes a general and flexible framework for pursuing discovery from unstructured data in a statistically principled way. The framework leverages recent methods from the literature on machine learning interpretability to map unstructured data points to high-dimensional, sparse, and interpretable “dictionaries” of concepts; computes statistics of dictionary entries for testing relevant concept-level hypotheses; performs selective inference on these hypotheses using algorithms validated by new results in high-dimensional central limit theory, producing a selected set (“discoveries”); and both generates and evaluates human-interpretable natural language descriptions of these discoveries. The proposed framework has few researcher degrees of freedom, is fully replicable, and is cheap to implement—both in terms of financial cost and researcher time. Applications to recent descriptive and causal analyses of unstructured data in empirical economics are explored. An open source Jupyter notebook is provided for researchers to implement the framework in their own projects.
@@ 1 Introduction
Empowered by recent developments in machine learning and AI, researchers in the social sciences are increasingly leveraging sources of unstructured data—such as text, audio, images, and videos—in quantitative analyses. In economics, interest in unstructured data sources is especially widespread, ranging from using speech recordings from FOMC meetings to better understand monetary policy [Gorodnichenko et al., 2023]; to using videos of start-up pitches to study entrepreneurship and investment [Hu and Ma, 2025]; to using open-ended survey questions to probe the economic behavior and beliefs of individuals [Haaland et al., 2024]; to using mug shots to study judicial decisionmaking [Ludwig and Mullainathan, 2024]; to using visual art or written narratives to infer long-run living standards [Gorin et al., 2025, Lagakos et al., 2025]; to using qualitative interview transcripts to better understand the impacts of RCT treatments [Bergman et al., 2024, Krause et al., 2025].
@@
In the social sciences, the promise of these unstructured data, fundamentally, is that they offer new opportunities to measure social phenomena that were previously unmeasurable. This includes both measures that a researcher is capable of pre-specifying before seeing the data as well as those that a researcher is not, and instead would like to “discover” through the data itself. When modern AI/ML methods are brought to bear on these goals, the former typically falls under the heading of “supervised learning,” whereby the researcher makes predictions of their ex-ante known quantity of interest using the unstructured data as features, and the latter typically falls under the heading of “unsupervised learning,” which is less amenable to a prediction-based framework, and demands that AI/ML methods uncover latent structure in the unstructured data with minimal researcher input. Though new statistical and econometric frameworks have recently become available for interpretable and statistically principled analyses of unstructured data in the supervised learning setting (e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025]), it is an open question as to how one best performs interpretable and statistically principled analyses of unstructured data in unsupervised learning settings that emphasize discovery. The framework proposed in this paper provides one possible answer to this open question.
Consider two motivating examples for the proposed framework, one focused on descriptive analysis, and another focused on causal inference. Both consider text as the unstructured data modality of interest, though the proposed framework applies more generally to other data modalities (see, e.g., the interpretability literature review in Section 2).
Example 1 (Information treatment RCT (c.f., Bursztyn et al. [2023])).
A researcher is interested in running a randomized controlled trial (RCT) with an information treatment. A key outcome of the experiment is the participant’s subjective reaction to the information treatment, and the researcher would like to discover any systematic differences in reactions across the treatment and control groups in an unsupervised way. As such, for all participants, the researcher collects text responses to open-ended questions eliciting reactions to the experimental intervention, and is interested in learning any relevant causal effects on the presence of concepts surfaced in these text responses.
Example 2 (Open-ended survey analysis (c.f., Stantcheva [2024])).
A researcher is interested in understanding the opinions and attitudes of some population of interest towards some particular (economic) policy. To learn about population attitudes without priming respondents to discuss particular topics, the researcher runs a survey on a representative sample from the population of interest that asks many open-ended questions about attitudes and beliefs, to which survey respondents reply in text. The researcher is interested in describing any prominent concepts that are surfaced in this open-ended survey.
@@
To achieve these motivating research goals—among many others—this paper proposes a general framework for conducting interpretable and rigorous statistical inference on unstructured data in the following way:
@@
    The researcher begins with an unstructured dataset they are interested in analyzing. Each unstructured data point is mapped to a “feature dictionary,” which is a binary vector; each element of the vector indicates the presence (or lack thereof) of a single, human-interpretable (“monosemantic”) feature in that unstructured data point. [^1] For more discussion of what “monosemanticity” means and its importance for LLM interpretability, see, e.g., the discussion in Bricken et al. [2023]. The creation of these interpretable feature dictionaries is accomplished using dictionary learning models newly popular in the machine learning interpretability literature for deep neural networks (DNNs) such as large language models (LLMs). These feature dictionary vectors are high-dimensional, and catalog a vast number of concepts. Intuitively, the feature dictionary should thought of as an inventory of all possible (monosemantic) concepts that, e.g., a LLM needed to learn for the purpose of next token prediction over a massive internet-scale corpus. By construction, these feature dictionary vectors are also sparse, i.e., only the few most “meaningful” concepts represented in the unstructured data point are cataloged as present by the dictionary. [^2] Meaningfulness, in a technical sense, relates to the reconstruction loss associated with learning the dictionary feature under a sparsity penalty; for more on this, see, e.g., Bricken et al. [2023].
    Example 1 (Continued).
    Each experimental participant’s post-treatment text response is passed through a LLM equipped with pretrained sparse autoencoder (SAE) [Bricken et al., 2023] (a popular implementation of a dictionary learning model); the SAE’s feature activations are pooled and thresholded to serve as a (binary) feature dictionary. There is one feature dictionary per text response (and one text response per experimental participant), each of which is of dimension on the order $p\approx 10^{5}$, though there are only on the order of $n\approx 10^{2}$ participants across both treatment and control groups. Feature number 1 in the dictionary may indicate the presence of sentences ending with exclamation points, feature number 2 may indicate the discussion of politics, and so on. [^3] For more examples of feature descriptions, see, e.g., https://www.neuronpedia.org/gemma-scope.
@@
    The researcher, now with access to a new dataset composed of feature dictionaries, formulates $p$ feature-specific hypothesis tests and associated test statistics—one for each concept cataloged in the feature dictionary. The researcher has great flexibility in designing relevant, feature-level hypotheses and t-stats at this stage of the framework, which may employ data sources beyond the feature dictionaries themselves.
    Example 1 (Continued).
    The researcher computes a difference in means across treatment and control groups for each feature indicator (binary entry in the feature dictionary) to test a total of $p$ null hypotheses of zero average treatment effect (ATE) on the probability of a given feature appearing in a text response.
@@
    All $p$ feature-specific hypotheses are tested using statistical procedures validated by novel results in high-dimensional multiple hypothesis testing and central limit theory. In particular, these procedures are shown to offer a novel formal guarantee over a generalized error rate for the selected set (the set of rejected null hypotheses) in high-dimensions, namely the $k$ familywise error rate ($k$-FWER) [Romano and Wolf, 2007]—the probability of making $k$ or more false rejections. This form of selective error control is less conservative that simple FWER control, [^4] Note that FWER is equivalent to $k$-FWER when $k=1$. leading to more powerful tests that permit the researcher to make more discoveries (rejections of true nulls). Moreover, the researcher may invert these tests to form generalized simultaneous confidence intervals, i.e., the probability that $k$ or more parameters are not contained by the random intervals is less than some desired level. These new formal results make only very mild assumptions on the data generating process, and may also be of independent interest in other high-dimensional selective inference settings.
    Example 1 (Continued).
    A fast, bootstrap-based algorithm is deployed to produce a critical value (or sequence of critical values) for rejecting each of the $p$ null hypotheses of no average treatment effect, guaranteeing $k$-FWER control in the selected set at level $\alpha$ (in large samples) for a researcher chosen value of $k$, even though $p\gg n$. These bootstrapped critical values may also be used to form generalized confidence intervals for the difference-in-means point estimates.
  4. 4.
    The output of the framework thus far is a selected set of rejected, feature-level null hypotheses (and corresponding generalized simultaneous confidence intervals and point estimates), a.k.a., “discoveries.” By virtue of the technology used to create them, however, the concepts encoded by the feature dictionary are only indexed numerically, and do not come with natural language descriptions—we know, e.g., that feature 1 likely encodes a human-interpretable concept, and that we rejected the null hypothesis associated with it, but we do not yet have an English description for what that concept is. As such, in the final stage of the framework, the researcher generates natural language descriptions of the concepts in the feature dictionary using a variant of a popular automatic interpretation (or “autointerp”) method that is localized to the unstructured dataset being explored. In order to rigorously evaluate the quality of these autointerp descriptions, we propose a new statistical formalization of a workhorse autointerp evaluation method, along with new estimators that permit statistically and conceptually principled autointerp quality evaluation.
    Example 1 (Continued).
    The researcher now has in hand some subset of the $p$ null hypotheses that have been rejected: a set of indices associated with particular elements of the feature dictionary. In order to provide a natural language description for each of the selected features for which there is evidence of a nonzero ATE (which are otherwise just identified by their numeric index), a dedicated “explainer LLM” reads examples of texts in which a given feature is present (or presents with high SAE activations) and offers an English interpretation for what concept is encoded by the feature; the credibility of this LLM-derived description of the feature is then evaluated with a held-out evaluation sample using newly proposed estimators.
  5. 5.
    The final output of the proposed framework is therefore: a set of feature-level discoveries (rejections of nulls) that come with generalized familywise error rate guarantees; natural (e.g., English) language descriptions of those discoveries; and numeric scores indicating the quality of those descriptions.
    Example 1 (Continued).
    The researcher discovers that the null of no average treatment effect is rejected for features indexed $\{4,101,5030,$...\}, meaning that with high probability the information treatment induces nonzero causal effects on the probability of the appearance of these features in texts collected across the treatment and control groups. Feature 4 is described by an explainer LLM as activating on “mentions of politics” and this description is evaluated to have a high “quality” score, i.e., the information treatment credibly causes increased discussion of politics; and so on.
@@
This particular framework for discovery from unstructured data conveys several scientific benefits. The first is that of having few researcher degrees of freedom. In many settings featuring exploratory analysis of unstructured data, there is concern that researchers may cherry pick aspects of their data to measure in order to reverse engineer a specific conclusion. By design, however, the framework proposed by this paper has very little room for motivated data snooping, as it requires no “human-in-the-loop” to direct what quantities are ultimately measured or how those quantities are interpreted. Off-the-shelf, open-source dictionary learning models generate $p$ features spanning a huge number of concepts, making it implausible that the researcher needs to define additional concepts to measure ad hoc, and making a failure to report hypotheses for certain features conspicuous. Localized autointerp methods, especially when coupled with formal interpretation quality evaluation, further limit a researcher’s ability to guide analysis towards a preordained conclusion. Moreover, prior to analysis, a researcher can easily preregister a choice of dictionary learning model and explainer LLM, hampering the possibility that a researcher applies this framework with many different models to select on favorable results. [^5] Even such a method for pursuing motivated data snooping is unlikely to succeed, given that, e.g., SAE features tend to exhibit some degree of universality across architectures (c.f., Wang et al. [2025]). The proposed framework is also inexpensive both in terms of financial cost and researcher burden; both empirical examples in this paper were computed on Google Colab notebooks using a single A100 GPU, access to which only costs tens of dollars per month (and is free for students). [^6] This pricing description is valid as of December 2025. This framework is easy to implement, requiring no human-in-the-loop to label unstructured data points, no additional AI/ML model training, and no additional data sources not already possessed by the researcher; the framework is nearly automatic in execution, and takes on the order of a single hour to run using the above-mentioned hardware (and with access to LLMs via API, it may only take minutes). This automaticity also lends the framework to fast and easy replication and sensitivity analysis, other important scientific benefits.
@@
    Novel, general-purpose results in high-dimensional central limit theory and high-dimensional multiple hypothesis testing that enable statistical procedures for controlling $k$-FWER in high-dimensional settings.
  3. 3.
    A new conceptual and statistical formalization of a popular framework for autointerp scoring, including novel estimators for implementing it.
The remainder of this paper is structured as follows: Section 2 discusses related literature; Section 3 describes the framework in detail (applied researchers are primarily encouraged to read Sections 3.1, 3.4, and 3.5); Section 4 applies the framework to two recent papers making use of unstructured data for discovery [Bursztyn et al., 2023, Stantcheva, 2024]; and Section 5 concludes.
@@ 2 Related Literature
Motivated by the capabilities of modern AI/ML methods for learning from unstructured economic data (see, e.g., Mullainathan and Spiess [2017], Gentzkow et al. [2019a], Ash and Hansen [2023], Dell [2025]), new econometric and statistical frameworks have been developed to facilitate principled statistical inference on low-dimensional features (predictions) learned from unstructured datasets under supervision, e.g., Angelopoulos et al. [2023], Ludwig et al. [2024], Carlson and Dell [2025], Rambachan et al. [2024]. This most recent wave of econometric literature emphasizes nonparametric frameworks compatible with black-box AI models coupled with debiasing methods, as opposed to model-based approaches, e.g., Gentzkow et al. [2019b], Battaglia et al. [2024]. Though the framework presented in this paper is concerned with principled inference on unstructured data without making parametric assumptions, it differs from this existing econometric literature in that its goal is primarily unsupervised discovery, as opposed to supervised detection, using AI/ML models.
Both the proposed framework and the recent econometric framework of Modarressi et al. [2025] support making causal inferences on text, with an emphasis on discovery. Modarressi et al. [2025] uses LLM reasoning as a tool for dimensionality reduction on the space of concepts present in text (i.e., as a low-dimensional hypothesis generation methodology) and incorporates human validation of LLM-selected hypotheses. In contrast, the framework proposed in this paper leverages recent advances in interpretability methods for LLMs in order to directly analyze the high-dimensional space of concepts present in text, and uses state-of-the-art autointerpretation methods to describe discovered concepts without further researcher intervention, similarly evaluating these descriptions using quantitative methods that require no researcher discretion. As such, these frameworks offer different profiles of scientific benefits and researcher costs for the problem of causal inference on text.
Modarressi et al. [2025] also directly leverages the work of Ludwig et al. [2017] in their pipeline, which is similarly relevant to the proposed framework, in that both emphasize discovery from high-dimensional data, and can likewise both be applied to the problem of causal inference on unstructured data (as it is used in Modarressi et al. [2025]). However, the proposed framework tackles the high-dimensional discovery problem using a novel high-dimensional selective inference procedure that maintains a high degree of interpretability, whereas Ludwig et al. [2017] modify the target of discovery to improve power, trading off interpretability to do so.
@@
There is an extensive literature on multiple hypothesis testing (or “selective inference” more broadly) in statistics, biostatistics, and economics (see, e.g., Romano et al. [2010] for a review in econometrics). The methods in this literature span asymptotic and finite sample valid frameworks, and low- and high-dimensional settings. In particular, for the problem considered in this paper, the literature on asymptotically valid, high-dimensional selective inference is most relevant, as such methods permit making very few assumptions about the data (c.f., assuming independence or PRDS of p-values in the case of Benjamini and Hochberg [1995]) and readily incorporate resampling methods that improve power (relative to guarding against worst-case dependence, c.f., Bonferroni corrections or the method of Benjamini and Yekutieli [2001]). The new results for high-dimensional multiple hypothesis testing appearing in this paper are derived by extending recent work on high-dimensional central limit theory for order statistics [Chernozhuokov et al., 2022, Ding et al., 2025]. The formal results presented in this paper are the first stated of their kind that enable asymptotic high-dimensional control of the $k$-FWER for test statistics that may be arbitrarily correlated with one another, complementing existing methods for asymptotic high-dimensional control of the FWER [Belloni et al., 2018] and the FDR under weak dependence [Liu and Shao, 2014, Belloni et al., 2018].
@@
There is a growing literature in economics and computer science on “hypothesis generation” at the pre-scientific stage of empirical inquiry [Ludwig and Mullainathan, 2024]. From the perspective of this literature, the proposed framework may be viewed as leveraging dictionary learning as a high-dimensional hypothesis generator, on top of which discovery is performed at the “scientific” stage of inquiry using high-dimensional multiple hypothesis testing procedures. As such, the framework in this paper contributes to the literature on hypothesis generation by providing a statistically principled pathway from hypothesis generation to hypothesis evaluation given access to only one (potentially small) unstructured dataset—a pathway that involves little to no researcher discretion about what generated hypotheses should be tested. The recent framework of Movva et al. [2025] uses dictionary learned features (specifically, sparse autoencoder features) explicitly for the purposes of hypothesis generation, further motivating the use of dictionary learning methods as tools for hypothesis generation in the social sciences, both in the present framework and beyond.
@@
There is a large literature in computer science on machine learning interpretability (see, e.g., Doshi-Velez and Kim [2017]). The machine learning interpretability methods leveraged by this framework originate from a nascent though highly active literature known as “mechanistic interpretability,” which seeks to develop methods for interpreting LLM behavior via quantitative analyses of model internals (e.g., activations, weights). In particular, the dictionary learning methods for DNN interpretability implemented in this framework were developed in Bricken et al. [2023], Templeton et al. [2024], i.e., sparse autoencoders (SAEs). SAEs are very wide autoencoders inserted at various layers of an LLM that are trained to intercept and reconstruct activations from model internals (such as the residual stream, MLPs or the like) under a sparsity-inducing penalization on a reconstruction loss. Under the “linear representation hypothesis” and “superposition hypothesis” (see, e.g., Bricken et al. [2023] for more discussion), these sparse autoencoders are thought to act as an overcomplete basis of the space of concepts in text leveraged by a LLM to make next token predictions, encouraging learned autoencoder latents to have monosemantic interpretations. There has been much debate about the success of SAEs (e.g., Leask et al. [2025]), though most criticisms still support the notion that SAEs and other dictionary learning methods have a comparative advantage in discovering as opposed to detecting concepts of interest in text [Peng et al., 2025]. Other dictionary learning methods that have been proposed as competitors to SAEs (e.g., transcoders [Paulo et al., 2025]) are similarly compatible with the framework proposed in this paper.
SAEs and other dictionary learning methods generate features, but not natural language feature descriptions. Based on the popularity of these methods, a related literature on “automatic interpretability” (or “autointerp”) methods has become active, which seeks to coherently use LLMs to describe the features discovered by dictionary learning methods at scale. Recent important papers in this literature include: Bills et al. [2023], Shaham et al. [2024], Paulo et al. [2024], Rajamanoharan et al. [2024]. Specifically, the autointerp framework proposed in this paper adapts and formalizes insights from the “detection scoring” method of Paulo et al. [2024] in a way that emphasizes “locality”: maintaining a high degree of relevance to the unstructured data generating process being considered.
Though the empirical examples considered in this paper handle one of the most popular unstructured data types—text—for which dictionary learning-based interpretability methods are the most mature, dictionary learning techniques have been successfully applied to many other modalities, including audio and images [Abdulaal et al., 2024, Bhalla et al., 2024, Fry, 2024, Daujotas, 2024, Pluth et al., 2025]. Future work would seek to apply the proposed framework to unstructured datasets in these other data modalities.